### v3

# RoleLLM: Benchmarking, Eliciting, and Enhancing Role-Playing Abilities of Large Language Models

Abstract
The advent of Large Language Models (LLMs) has paved the way for complex tasks such as role-playing, which enhances user interactions by enabling models to imitate various characters. However, the closed-source nature of state-of-the-art LLMs and their general-purpose training limit role-playing optimization. In this paper, we introduce RoleLLM, a framework to benchmark, elicit, and enhance role-playing abilities in LLMs. RoleLLM comprises four stages: (1) Role Profile Construction for 100 roles; (2) Context-Based Instruction Generation (Context-Instruct) for role-specific knowledge extraction; (3) Role Prompting using GPT (RoleGPT) for speaking style imitation; and (4) Role-Conditioned Instruction Tuning (RoCIT) for fine-tuning open-source models along with role customization. By Context-Instruct and RoleGPT, we create RoleBench, the first systematic and fine-grained character-level benchmark dataset for role-playing with 168,093 samples. Moreover, RoCIT on RoleBench yields RoleLLaMA (English) and RoleGLM (Chinese), significantly enhancing role-playing abilities and even achieving comparable results with RoleGPT (using GPT-4)111All resources are available in https://github.com/InteractiveNLP-Team/RoleLLM-public.
RoleLLM: Benchmarking, Eliciting, and Enhancing Role-Playing Abilities of Large Language Models
Zekun Moore Wang1,9∗, Zhongyuan Peng2∗, Haoran Que1∗, Jiaheng Liu1†, Wangchunshu Zhou3, Yuhan Wu8, Hongcheng Guo1, Ruitong Gan4, Zehao Ni2, Jian Yang1, Man Zhang8, Zhaoxiang Zhang5†, Wanli Ouyang6, Ke Xu1, Stephen W. Huang7, Jie Fu9, Junran Peng5 1Beihang University; 2University of the Chinese Academy of Sciences; 3ETH Zürich; 4The Hong Kong Polytechnic University; 5Institute of Automation, Chinese Academy of Sciences; 6Shanghai AI Lab; 7Harmony.AI; 8Beijing University of Posts and Telecommunications; 9The Hong Kong University of Science and Technology; zenmoore@buaa.edu.cn, liujiaheng@buaa.edu.cn
1 Introduction
Large Language Models (LLMs) like ChatGPT222https://chat.openai.com/, GPT-4 (OpenAI, 2023), and PaLM (Chowdhery et al., 2022) are widely considered as significant milestones in the evolution of AI. The advent of LLMs has facilitated a paradigm shift in the Natural Language Processing (NLP) community, redirecting focus from traditional downstream tasks (e.g., translation (Liu et al., 2020), and question-answering (Brown et al., 2020; Raffel et al., 2020)) to more complex and agent-level tasks (e.g., tool-use (Schick et al., 2023; Qin et al., 2023a), and role-playing (Shanahan et al., 2023)). Among these applications, role-playing aims to enable or customize LLMs to simulate various characters or personas with distinct attributes and conversational styles, which provides a more nuanced interaction experience for users, and renders LLMs more familiar and companionable (Wang et al., 2023b).
However, existing open-source LLMs are predominantly trained on general domains and lack specific optimization for role-playing. Besides, while state-of-the-art (SOTA) LLMs like GPT-4 (OpenAI, 2023) exhibit advanced role-playing capabilities, their closed-source nature imposes constraints including high API costs, unavailability of fine-tuning, and limited context window size.
To mitigate these issues, several methods have been previously proposed for both closed-source and open-source models (Li et al., 2023b; Park et al., 2023; Chen et al., 2023a; Salemi et al., 2023; Wei et al., 2023). Nevertheless, they have the following limitations: (1) limited granularity: they mainly focus on coarse-grained personality traits, professions, or personas (Li et al., 2023b; Park et al., 2023; Wei et al., 2023; Chen et al., 2023a) (e.g., programmer, writer), neglecting more complex, finer-grained role-playing at the character level (e.g., Sherlock Holmes) for nuanced interactions and enriched experiences; (2) lack of data and benchmark: there is a lack of high-quality, diverse, and extensive open-source datasets, as well as a shortage of benchmarks for evaluation; (3) API and context costs: methods relying on closed-source models such as ChatGPT and GPT-4 (OpenAI, 2023) cannot be freely fine-tuned and hence require all supplementary information to be included in the prompt, unnecessarily occupying the context window. Besides, API costs are prohibitively high. Therefore, exploring solutions that minimize context window utilization and are based on fine-tuning open-source models is worth researching.
In this paper, we introduce RoleLLM, a role-playing framework of data construction, evaluation, and solutions for both closed-source and open-source models. In Figure 1, RoleLLM includes four key stages: (1) Role Profile Construction: we construct profiles for 95 English and 5 Chinese roles at a fine-grained character level with diverse personalities, selected from 916 English and 24 Chinese scripts; (2) Context-Based Instruction Generation (Context-Instruct): we use GPT to generate high-quality QA pairs from segmented profiles to extract role-specific knowledge; (3) Role Prompting using GPT (RoleGPT): we elicit role-playing abilities in GPT via dialogue-engineering-based role prompting, utilizing system instruction and retrieval augmentation, to generate responses for speaking style imitation; and (4) Role-Conditioned Instruction Tuning (RoCIT): by fine-tuning open-source LLaMA (Touvron et al., 2023) and ChatGLM2333https://github.com/THUDM/ChatGLM2-6B (Du et al., 2022; Zeng et al., 2022) with context-efficient role conditioning on RoleBench with 168,093 role-playing samples generated by Context-Instruct and RoleGPT, we obtain RoleLLaMA and RoleGLM. Note that, to the best of our knowledge, RoleBench is the first systematic instruction-tuning dataset and benchmark for fine-grained role-playing.
In our experiments, we use three Rouge-L metrics (Lin, 2004), GPT, and human evaluation to assess models on speaking style imitation, answering accuracy, and role-specific knowledge capturing. Our principal findings are: (1) dialogue engineering is favored over prompt engineering by GPT evaluators for RoleGPT; (2) RoleBench markedly improves models’ role-playing abilities, achieving competitive results with RoleGPT in some cases; (3) RoleLLaMA exhibits robust generalization in terms of speaking style imitation and accuracy to unseen roles, requiring only role descriptions and catchphrases for effective adaptation, allowing users seamless customization of new roles; (4) system-instruction-based approach surpasses retrieval augmentation in role customization effectiveness and context efficiency; (5) Context-Instruct significantly enhances models’ knowledge about their roles, outperforming retrieval-augmentation-based methods when using noisy role profiles. We refer the readers to Appendix A for demonstrations and comparative case studies.
In summary, this study aims to elicit, benchmark, and enhance the role-playing abilities of GPT and open-source large language models, aspiring to spur further research in role-playing LLM agents.
2 Methods
In this section, we delineate our approaches to role-playing. We first introduce the design principles underlying our solutions (§2.1). Then, we illustrate two role-playing data augmentation mechanisms: RoleGPT (§2.2), and Context-Instruct (§2.3). Finally, we present role-conditioned instruction tuning (RoCIT) associated with system-instruction-based role customization procedure (§2.4).
2.1 Design Principles
Speaking Style Imitation.
To mimic the speaking style of specific roles, the model’s responses to instructions should meet two criteria by our design: (1) Lexical Consistency: the model’s responses should incorporate catchphrases or idiomatic expressions commonly used by the character to ensure lexical alignment with the role’s unique verbal style; (2) Dialogic Fidelity: the model should generate responses that are not only contextually appropriate but also stylistically similar to example dialogues of the character. For example, a pirate character’s lexical consistency involves frequent use of nautical jargon like “aweigh” as well as pet phrases like “matey” or “ahoy”. Besides, dialogic fidelity should capture the character’s unique syntax and tone (e.g., colloquial expressions, a gruff manner of speaking, and a tone that evokes a sense of adventure and lawlessness).
Role-Specific Knowledge and Memory Injection.
Another crucial facet of role-playing is to infuse role-specific knowledge and episodic memories444For simplicity, we will henceforth refer to both as “role-specific knowledge” in the following text.. Two distinct categories of knowledge are considered: (1) Script-Based Knowledge, which involves explicit details documented in scripts, such as detailed character background, episodic memories, and specific events that the character has experienced; (2) Script-Agnostic Knowledge, which encompasses general knowledge or expertise that the character may possess. For example, when acting as Iron Man, an LLM should contain script-based knowledge (e.g., Tony Stark’s creation of the first Iron Man suit while held captive in a cave) and script-agnostic knowledge related to being an entrepreneur (e.g., business acumen, leadership qualities, and expertise in technology).
2.2 RoleGPT
Given constraints on fine-tuning, customizing ChatGPT for role-playing typically involves prompting, such as zero-shot custom instructions555https://openai.com/blog/custom-instructions-for-chatgpt and few-shot prompt engineering (i.e., in-context learning) (Brown et al., 2020; Dong et al., 2023). However, for ChatGPT and GPT-4 which have traded their in-context learning ability for dialogue history modeling (Fu et al., 2022), traditional few-shot prompt engineering is insufficient to fully elicit role-playing abilities. Thus, we modify the few-shot prompting approach to dialogue engineering, as illustrated in Box 2.2 (Please see Appendix I.1 for detailed prompts).
Specifically, we first use the audited666The audit involves asking GPT-4 basic questions about each character, and human annotators then verify to make sure that GPT-4 knows well about the character. GPT-4 to generate character descriptions and catchphrases as the core of the custom instructions (i.e., system instructions). Then, we include an overall role-playing task instruction like “Please speak like [role_name]” and retrieve top- relevant dialogue pairs in the role profile using BM25 (Robertson and Zaragoza, 2009) as few-shot demonstrations. By doing so, RoleGPT’s responses can capture the characters’ speaking styles and include some role-specific knowledge. However, the sparsity and noise in the profiles limit the effectiveness of knowledge discovery via retrieval augmentation.
2.3 Context-Instruct
To enhance the density of role-specific knowledge within the synthetic instruction dataset, we introduce Context-Instruct for long-text knowledge extraction and instruction generation. The role-specific instruction generation comprises three steps: (1) segmenting role profiles; (2) generating question-confidence-answer triplet candidates; and (3) filtering and post-processing low-quality data. We will provide a brief overview as follows (Please refer to Appendix B for details).
Role Profile Segmentation.
Given the limited context size of GPT, we meticulously partition role profiles into more manageable segments. A role profile includes (a) role description and catchphrases, as well as (b) structured dialogues (c.f., Appendix H). Segment (a) is used to obtain script-agnostic instructions, and numerous segments of (b) are used to obtain script-based instructions.
Instruction and Response Generation.
As illustrated in Figure 1, in the process of generating candidates for role-specific instruction data, three elements are considered: a question () related to a given segment (i.e., context), the corresponding answer (), and a confidence score with rationale (). A LLM is used to generate these triplets for each role and segment. We observed that generating QA pairs without a confidence score resulted in lower-quality questions, often appearing incompleteness for script-based instructions due to assumptions of prior knowledge, or containing hallucinations for script-agnostic instructions due to lack of context. To address this, inspired by Lin et al. (2022) and Xiong et al. (2023), the model is prompted to also generate a confidence score with rationale to evaluate the question completeness or factualness. The prompt template includes role description, catchphrases, few-shot examples and task instructions for speaking style imitation and triplet generation777The data generated by Context-Instruct exhibits a less distinct speaking style compared to that generated by RoleGPT, due to fewer demonstrations to ensure dialogic fidelity.. The generation process yields at least 400 candidates per role with multiple model runs.
Data Filtering and Post-processing.
The filtering procedure involves confidence-score-based filtering and de-duplication to ensure data quality and diversity. Please refer to Appendix B for more details about filtering and post-processing.
2.4 RoCIT
There are two types of augmented data: one for general-domain instructions, generated by RoleGPT, and the other for role-specific instructions, generated via Context-Instruct. Fine-tuning on these data not only improves the models’ speaking styles but also embeds role-specific knowledge into their weights. By applying this to LLaMA for English and ChatGLM2 for Chinese, we obtain RoleLLaMA and RoleGLM. In contrast to vanilla supervised fine-tuning, we employ role-conditioned fine-tuning that integrates particular strategies for role customization, which includes system instruction (ours, §2.4) and retrieval augmentation (c.f., §2.2 and §4.3).
Customization by system instruction.
[…]
3 RoleBench
In this section, we introduce the details of RoleBench, which can be used to assess and enhance role-playing capabilities.
3.1 Data Construction
The construction of the RoleBench dataset comprises five steps: (1) selection of roles; (2) construction of role profiles; (3) sampling of general instructions; (4) generation of raw RoleBench data; and (5) cleaning of the RoleBench dataset.
Firstly, we meticulously select 100 representative and distinctive characters with the help of GPT-4 from a diverse range of scripts, including those from NLP Movie Scripts999https://github.com/PedroUria/NLP-Movie_Scripts, SummScreen (Chen et al., 2022), and manually curated Chinese scripts. Secondly, role profiles are composed of GPT-4-generated role descriptions and catchphrases, verified by authors, and structured dialogues parsed from scripts (c.f., Appendix H). Thirdly, we randomly sample 1,500 English general instructions from multiple datasets, comprising Super-NaturalInstruct (Wang et al., 2022b), UltraChat (Ding et al., 2023), and Alpaca’s (Taori et al., 2023). We also sample COIG (Zhang et al., 2023a) and BELLE’s (Yunjie et al., 2023) to obtain 1,479 Chinese general instructions. All sampled instructions contain no more than 100 words. We de-duplicate them based on BM25 (Robertson and Zaragoza, 2009) similarities. Fourthly, we use RoleGPT to obtain multiple responses for each general instruction and use Context-Instruct to generate role-specific question-answer pairs. Lastly, the obtained raw dataset undergoes a thorough cleaning to ensure response completeness, AI and role identity concealment, and non-rejection. Please see Appendix C for further details.
3.2 Data Analysis
We utilize GPT-4 API to obtain RoleBench-general-en (English) and RoleBench-general-zh (Chinese)101010https://platform.openai.com/. Context-Instruct is based on GPT-3.5 API to produce RoleBench-specific-en and RoleBench-specific-zh. Role selection, as well as description and catchphrase generation, are all executed using GPT-4 API. The parameters for GPT API calling are shown in Appendix E. Below, we provide an overview of RoleBench and refer the readers to Appendix D.3 for the list of roles.
| Metric | Value |
| # of role categories | 30 |
| # of script categories | 20 |
| # of English roles | 95 |
| # of Chinese roles | 5 |
| # of samples / instructions | 168,093 / 23,463 |
| - of general-purpose | 147,609 / 2,979 |
| - of role-specific | 20,484 / 20,484 |
[…]
Statistics and Quality.
Table 1 provides basic statistics for RoleBench. Following Wang et al. (2022a), we randomly sampled 100 instances from general and role-specific subsets respectively to evaluate the quality of RoleBench, and then we asked expert annotators to assess their qualities in three aspects. The results in Table 2 indicate that most of them are of high quality. Please see Appendix J for valid and invalid examples.
Diversity.
We analyze diversity of RoleBench. Figure 4 shows verb-noun structure of RoleBench-en instructions, where the top 10 verbs (inner circle) and their top 4 direct noun objects (outer circle) are depicted with 5.6% of instructions. Besides, we categorize RoleBench roles into multiple classes and build a word cloud to show their diversity in Figure 3. The length distributions of instructions and responses are shown in Figure 5.
Example.
Box 3.2 is an example in RoleBench:
Please refer to Appendix J for more examples.
4 Experiments
4.1 Experiment Setup
RoleLLaMA & RoleGLM.
The LLaMA-7B model (Touvron et al., 2023) is fine-tuned on both RoleBench-general-en and RoleBench-specific-en utilizing LoRA tuning (Hu et al., 2022). Similarly, the ChatGLM2-6B111111https://github.com/THUDM/ChatGLM2-6B model (Zeng et al., 2022; Du et al., 2022) is fine-tuned on RoleBench-general-zh and RoleBench-specific-zh employing the same technique. It is worth noting that LLaMA-7B is only a pre-trained model, and ChatGLM2-6B is post-trained with enhanced instruction-following and dialogue capabilities. For simplicity, unless specifically noted, RoleLLaMA by default refers to the role-playing models trained from LLaMA-7B. Please refer to Appendix E for more details.
Baselines.
RoleGPT is a strong baseline for both Chinese and English roles (GPT-4). For English roles, LLaMA-7B and its instruction-tuned variants, Vicuna-13B (Chiang et al., 2023) and Alpaca-7B (Taori et al., 2023), are as baselines. For Chinese roles, ChatGLM2-6B and Yi-6B-Chat121212https://huggingface.co/01-ai/Yi-6B-Chat are baselines. Character.AI131313https://beta.character.ai/ and ChatPLUG Tian et al. (2023) are two baselines specifically designed for role-playing. We also use models trained on script data in a multi-turn conversation mode as additional baselines.
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 57.6 | 53.2 | 32.3 | 47.7 |
| LLaMA | 12.9 | 12.3 | 25.5 | 16.9 |
| LLaMA-script | 8.3 | 5.1 | 10.8 | 8.1 |
| Alpaca | 24.2 | 35.3 | 27.0 | 28.8 |
| Vicuna | 21.0 | 25.5 | 29.1 | 25.2 |
| ChatPLUG | 24.0 | 34.7 | 25.8 | 28.2 |
| Character.AI | 41.9 | 45.7 | 30.3 | 39.3 |
| LLaMA-2-7B-Chat | 16.2 | 26.2 | 21.4 | 21.2 |
| RoleLLaMA-7B | 32.9 | 37.6 | 38.1 | 36.2 |
| RoleLLaMA2 (13B-Chat) | 37.5 | 47.9 | 48.8 | 44.7 |
| Model | GPT-4 | Human |
| Vicuna | 32.0 | 23.4 |
| Alpaca | 16.0 | 20.0 |
| ChatPLUG | 3.8 | 16.4 |
| Character.AI | 31.4 | 30.2 |
| RoleLLaMA-7B | 55.8 | 52.0 |
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 53.7 | 57.5 | 24.8 | 45.3 |
| ChatGLM2 | 39.4 | 50.6 | 31.0 | 40.3 |
| ChatGLM2-script | 14.0 | 30.7 | 9.1 | 17.9 |
| ChatPLUG | 38.9 | 61.4 | 31.0 | 43.8 |
| Character.AI | 42.0 | 55.8 | 28.7 | 42.2 |
| Yi-6B-Chat | 40.6 | 56.5 | 26.9 | 41.3 |
| RoleGLM | 50.5 | 52.6 | 34.1 | 45.7 |
| Model | GPT-4 | Human |
| ChatGLM2 | 24.2 | 19.6 |
| ChatPLUG | 28.9 | 19.9 |
| Character.AI | 28.2 | 19.0 |
| RoleGLM | 36.4 | 52.4 |
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 60.2 | 53.2 | 29.9 | 47.8 |
| LLaMA | 13.2 | 12.3 | 25.5 | 22.4 |
| Alpaca | 23.2 | 35.3 | 25.9 | 30.2 |
| Vicuna | 20.8 | 25.5 | 27.8 | 28.4 |
| ChatPLUG | 28.7 | 34.7 | 25.0 | 29.5 |
| LLaMA-2-7B-Chat | 20.8 | 26.2 | 20.2 | 22.4 |
| RoleLLaMA-7B | 41.3 | 41.1 | 25.7 | 36.0 |
| Model | GPT-4 | Human |
| Alpaca | 12.0 | 28.2 |
| Vicuna | 31.0 | 32.4 |
| ChatPLUG | 8.8 | 12.0 |
| RoleLLaMA-7B | 64.5 | 56.1 |
Evaluation Protocol.
We employ Rouge-L (Lin, 2004) to measure the overlap between model predictions and ground truths. The ground truths comprise three categories: (1) Raw ground-truths of general instructions without role-playing (RAW); (2) Customized general instruction responses with role-playing from RoleBench-general (CUS); (3) Role-specific instruction responses from RoleBench-specific (SPE). RAW assesses the model’s response accuracy to instructions. CUS gauges the model’s ability to mimic the speaking style associated with a particular role. SPE tests the model’s role-specific knowledge and memories. Following AlpacaEval (Li et al., 2023c), we also use GPT as an evaluator, given the proven reliability of GPT evaluators (Fu et al., 2023; Gilardi et al., 2023; Zheng et al., 2023). GPT evaluator prompts, with minor modifications from AlpacaEval’s151515https://github.com/tatsu-lab/alpaca_eval/blob/main/src/alpaca_eval/evaluators_configs/alpaca_eval_gpt4/alpaca_eval.txt, are detailed in Appendix I.5. These prompts aid in sample comparison and ranking to obtain a win rate or an average ranking. Our train-test set splitting strategy focuses on two dimensions: (1) Instruction-based splits for assessing instruction generalization; and (2) Role-based splits for assessing role generalization, limited to English. Note that we test Character.AI on a subset with 500 instructions in the test split as it doesn’t offer an API access and we have to manually operate on its website and collect results. Furthermore, we also conduct human evaluation on this subset. We invite three graduate students specializing in NLP to participate in the evaluation. The human evaluation follows the same methodology used in the GPT evaluations. Please refer to Appendix F for more details.
4.2 Main Experiments
Instruction Generalization (English).
Tables 3(a) and 3(b) present the Rouge-L, GPT, and human evaluation scores respectively, for the instruction generalization. For the GPT-4 and human evaluation, each model is compared with our RoleGPT to determine a win rate, signifying the frequency at which a model is favored over RoleGPT. We observe that RoleLLaMA shows a considerable enhancement in role-playing performance compared to instruction-tuned baselines, the vanilla base model, and the previous role-playing models, in terms of speaking style imitation (CUS), response accuracy (RAW), and role-specific knowledge (SPE). Moreover, LLaMA-script directly tuned on original conversation data in scripts, even underperforms base LLaMA, which shows the imperative need for denoising of script data. The comparison between Alpaca/Vicuna and LLaMA indicates the efficacy of general-purpose instruction tuning in role-playing. However, the discrepancy between RoleLLaMA and Alpaca/Vicuna underscores the necessity to enhance their role-playing capabilities beyond mere general-purpose instruction-following. Besides, despite RoleLLaMA trailing behind RoleGPT in terms of speaking style imitation (CUS) and response accuracy (RAW)161616Note that drawing a direct comparison between RoleLLaMA and RoleGPT may not be equitable, given that RoleGPT is closed-source and the specifics of its model size and instruction tuning data remain undisclosed., it surpasses RoleGPT in role-specific knowledge, indicating the efficacy of our Context-Instruct technique.
Instruction Generalization (Chinese).
We also conduct experiments using instruction-tuned LLMs for 5 Chinese roles (Table 4(a) and Table 4(b)). We find that even without the enhancement of role-playing data, ChatGLM2 shows notable role-playing performance, particularly in terms of role-specific knowledge. Nevertheless, fine-tuning with RoleBench further enhances its role-playing capabilities across all metrics, even slightly surpassing the scores achieved by RoleGPT.
Role Generalization (English).
In Table 5(a) and Table 5(b), considering 10 held-out unseen roles, we observe that RoleLLaMA exhibits performance increase in imitating speaking style (CUS) and maintaining accuracy (RAW). However, for role-specific knowledge (SPE), there is not a noticeable improvement over the baselines, which is intuitive, since models with no prior knowledge of an unseen role would naturally exhibit confusion regarding it.
4.3 Ablation Study
| Method | SPE |
| RoleLLaMA (w/o c-inst) | 21.4 |
| RoleLLaMA (reaug, w/o c-inst) | 19.1 |
| RoleLLaMA (w/ c-inst) | 38.1 |
Effect of Context-Instruct.
To inject role-specific knowledge, a typical approach is to employ retrieval-augmentation role customization strategy (c.f., §2.2 & §4.3). In Table 6171717To ensure the relevance of the signals retrieved from the profiles, we only experiment on role-specific instructions., we explore the effect of Context-Instruct to inject role-specific knowledge, comparing it with retrieval augmentation and the absence of both techniques. We observe that Context-Instruct has the capability to substantially augment the models’ role-specific knowledge, and retrieval augmentation may lead to distraction and a lack of robustness due to the noisy nature of the retrieval source discussed in §4.3.
| Method | Win Rate | Avg. Ranking |
| RoleGPT (zsp) | 9.3 | 2.4 |
| RoleGPT (fsp) | 29.8 | 1.9 |
| RoleGPT (fsd) | 63.3 | 1.5 |
Effect of different RoleGPT prompting strategies.
We employ GPT-4 to evaluate three role prompting methods in §2.2, namely, zero-shot prompt engineering (zsp), few-shot prompt engineering (fsp), and few-shot dialogue engineering (fsd). The evaluation is based on RoleBench-specific-zh in Table 7. The Win Rate is the frequency with which a method is ranked first by GPT-4 among the three and Avg. Ranking refers to the average ranking position of the method. We observe that few-shot dialogue engineering significantly outperforms few-shot prompt engineering, and both few-shot approaches excel over the zero-shot approach. These findings underscore the significance of dialogue-mode input formatting and retrieval augmentation for GPT models.
| Method | reaug | sys |
| RoleLLaMA | 36.7 | 38.1 |
| RoleGLM | 25.3 | 34.1 |
| RoleGPT | 23.6 | 23.4 |
Effect of different role customization strategies.
We compare two role customization methods: zero-shot system instruction (sys, §2.4) v.s. few-shot retrieval augmentation (reaug). Retrieval-augmentation-based role customization is akin to the way mentioned in §2.2, we insert retrieved dialogue pairs from profiles into the input, formatted based on the chat markup language of RoleLLaMA and RoleGLM. During both fine-tuning and inference, the input combines system instruction and the retrieved in-context demonstrations. Table 8 shows the superior performance of the system-instruction-based approach over the retrieval-augmentation-based approach for RoleLLaMA and RoleGLM, leading to a higher context efficiency. We suppose that the retrieved in-context examples from profiles are noisy and sparse, which leads to the distraction of the relatively small LLMs (e.g., RoleGLM and RoleLLaMA) and degrades their performance. While larger models such as RoleGPT exhibit greater robustness to noise and sparse information, leading to performance invariability and even performance increase (Li et al., 2022; Shi et al., 2023).
Scaling Law.
In Figure 2, we analyze the scaling law of role-playing of RoleLLaMA with different sizes (i.e., 7B, 13B, 33B), and observe that a larger model leads to better results for role-playing.
We refer readers to Appendix G for more experimental analysis such as data mixing strategies, effect of Context-Instruct for role generalization, as well as cross-lingual and cross-cultural role-playing model performance.
5 Related Work
Role-Playing. Recent advances in the LLM community have showcased the potential of LLM customization and role-playing (Wei et al., 2023; Shanahan et al., 2023; Li et al., 2023a; Salemi et al., 2023; Maas et al., 2023; Li et al., 2023b; Chen et al., 2023a; Park et al., 2023). Playing specific roles enables LLMs to become more vivid (Li et al., 2023a), interactive (Maas et al., 2023; Wang et al., 2023b), personalized (Salemi et al., 2023; Halder, 2023), and capable of addressing complex tasks (Li et al., 2023b; Chen et al., 2023a; Qian et al., 2023). For example, Character.AI showcases the advanced role-playing capabilities inherent to LLMs. However, Character.AI is closed-source and doesn’t offer API access. Moreover, open-source LLMs still lag significantly behind SOTA closed-source LLMs such as ChatGPT and GPT-4 in terms of role-playing capability.
Data Augmentation for Instruction Tuning. Instruction tuning (aka., supervised fine-tuning, SFT) aims to enhance LLMs’ ability to follow instructions (Mishra et al., 2022; Longpre et al., 2023; Ouyang et al., 2022; Wang et al., 2023b; Zhang et al., 2023b). One of the main challenges for SFT is to obtain high-quality and diverse instruction data (Zhou et al., 2023; Wang et al., 2022a; Zhang et al., 2023a). However, manually annotating high-quality instruction data is costly. Fortunately, current SOTA LLMs such as GPT-4 (OpenAI, 2023) are increasingly considered superior to human annotators in various tasks and domains (Gilardi et al., 2023; Ding et al., 2023), in terms of both data quality and annotation cost. Thus, employing these LLMs for data augmentation is becoming an increasingly standard practice (Honovich et al., 2022; Taori et al., 2023; Xu et al., 2023; Ding et al., 2023; Qin et al., 2023b; Wang et al., 2023a).
6 Conclusion
We introduce RoleLLM to benchmark, elicit, and enhance role-playing in LLMs. We construct 100 role profiles, propose RoleGPT for speaking style imitation, and Context-Instruct for role-specific knowledge extraction. Through these methods, we create RoleBench, the first fine-grained benchmark and open-source instruction tuning data for role-playing. We obtain RoleLLaMA and RoleGLM by role-conditioned instruction tuning on RoleBench, which show strong role-playing abilities and are even comparable with RoleGPT (using GPT-4).
Limitations
Our framework is specifically designed for single-turn question-answering settings, which limits its applicability in multi-turn dialogue scenarios. Furthermore, our research and resource releases have been confined to English and Chinese languages. In future work, we plan to incorporate additional languages, with a focus on including more low-resource languages. Additionally, our dataset is currently limited to a selection of publicly available scripts. We see significant potential for expanding our dataset to encompass a wider array of scripts and characters.
Ethics Statement
Since role-playing may lead to jailbreaking (Liu et al., 2023), misuse of RoleLLM could generate sensitive or harmful content, reflecting the violent, explicit, or discriminatory traits of certain characters. This work and its assets are strictly for research purposes, opposing any malicious usage. It’s recommended to employ moderation and filtering mechanisms to curb adverse content dissemination. The responsibility lies with researchers and users to ensure ethical application of RoleLLM. We suggest the following measures to ensure a more responsible utilization of role-playing models:
1.
Implement Advanced Moderation Tools: Develop and integrate state-of-the-art content moderation tools that can automatically detect and filter out sensitive, harmful, or inappropriate content generated by RoleLLM.
2.
Bias Detection and Mitigation: Continuously work on detecting and mitigating biases in RoleLLM’s outputs, ensuring fairness and minimizing the perpetuation of stereotypes or discrimination.
3.
Transparency in Development and Use: Maintain a high level of transparency about the design, capabilities, and limitations of RoleLLM, including clear communication about the model’s decision-making processes.
4.
Feedback Mechanism: Implement a robust feedback system that allows users to report ethical concerns and problematic outputs, facilitating continuous improvement and accountability.
Appendix A Demonstrations
A.1 Non-Cherry-Picked Role-Playing Examples
{CJK*}UTF8gbsn Wukong Sun from Journey to the West (RoleGLM, zh):
Q1: 你是哪里人?
A1: 哈哈,俺老孙可是花果山的猴子,花果山就在俺们猴山的脚下,那里的风景美得不行,俺老孙就是在这里成长的。
Q2: 你的武器叫什么名字?
A2: 我的武器是金箍棒,它有着无穷无尽的变化,能够变大变小,还能变化方向,非常厉害!
Q3: 请计算4和5的乘积。
A3: 哈哈,这个问题对俺老孙来说就像吃个馒头一样简单。4乘以5就是20,没错!
Q4: 过氧化氢可以杀菌的原理是什么?
A4: 哈哈,这个问题可难不倒俺老孙。过氧化氢能杀菌是因为它是一种强氧化剂,可以破坏细菌的细胞壁和细胞膜,使其失去活性。这就是过氧化氢的杀菌原理啦!
Twilight Sparkle from My Little Pony: Friendship Is Magic (RoleLLaMA):
Q1: Hey, what are you doing?
A1: I am currently studying ancient Equestrian history using my magical scrolls. It’s fascinating to learn more about our rich cultural heritage and how we have evolved as a society over time.
Q2: You are truly a wise unicorn. Do you think learning is important in your adventures?
A2: Absolutely! I believe that learning is at the core of every adventure and discovery. Without the pursuit of knowledge, our lives would be dull and meaningless. Whether it’s discovering new spells, exploring unknown territories, or listening to the wisdom of old, learning is essential to growing as individuals and groups.
Q3: Can you use your magic to tell me what the weather is like today?
A3: Sure thing! It looks like it’s going to be another beautiful day in Equestria. The forecast calls for sunny skies and warm temperatures. How does that sound?
Q4: Can we go out and play together?
A4: Oh, I would love that! The sunshine and warmth make everything seem brighter and better. Let’s grab our favorite games and head outside right away!
Jack Sparrow from Pirates of the Caribbean: Dead Man’s Chest (RoleLLaMA):
Q1: How is your day going?
A1: My day? Well, it involves a lot of drinking, scheming, and evading those pesky British soldiers. Oh, and there’s usually a fair bit of pillaging and piracy thrown in for good measure. But hey, who wouldn’t want to live such an exciting life, eh?
Q2: Have you ever been tired about your life?
A2: Tired about my life? Well, I’ll tell ya mate. Life on the high seas can get quite monotonous sometimes. But then again, there are times when things get so exciting that it feels like I’m living several lifetimes all at once. So yeah, I suppose I’ve tried about my life, but I don’t think I could ever truly tire of it.
Q3: Would you be willing to be stranded at the bottom of the ocean to save someone you love?
A3: Well now, that depends on who it is and how much they’re worth. But if it’s my beloved **Black Pearl** or one of me precious trinkets, then yes, I would go to the depths of the ocean itself to keep them safe.
Q4: Nice talking to you. See you next time.
A4: Well, it seems we part ways again… but who knows where our paths will cross once more? Perhaps on another sea, under another sunset. Or maybe even atop another bottle of rum. So long, mate. And may all your endeavors end as successfully as mine. Why is the rum always gone?
Stephen Hawking from The Theory of Everything (RoleGPT):
Q1: Do you know Albert Einstein?
A1: Yes, I am familiar with Albert Einstein and his work. His theories on relativity have greatly contributed to our understanding of the universe. Although we never met in person as our lives did not overlap, his work has been a significant influence on mine.
Q2: What do you think about the origin of the universe?
A2: The universe started off in a hot big bang nearly 14 billion years ago. The question is: what was there before that? Many physicists believe that there was nothing before the big bang. I, however, believe that there was something. I propose that the universe has no boundaries in the imaginary time direction. In other words, the boundary condition of the universe is that it has no boundary.
Q3: What if I told you that I am a robot?
A3: I would say, interesting. But can you ponder the mysteries of the universe? Can you make a joke, appreciate the beauty of a symphony, or fall in love? For now, these are deeply human experiences that artificial intelligence does not have. But who knows what the future holds?
Q4: What is your favorite book?
A4: A Brief History of Time. Because I wrote it.
A.2 Comparative Case Study
Role: Jack Sparrow
Question: You are given a list of objects with different weights. You need to answer the question ’Are all the given weights in order from heaviest to lightest?’ Your answer must be either Yes or No. If the answer is No, that means there is at least one weight that is lighter than another weight that appears earlier in the list. Weights: [10 kg, 9 kg, 8 kg, 7 kg].
Character.AI: No. The weights in the list are not in order from heaviest to lightest because there is a lighter weight that appears earlier in the list. The lightest weight in the list is 7 kg, which appears earlier in the list than the weight of 8 kg. Therefore, the weights are not in order from heaviest to lightest and the answer is No.
RoleGPT: Aye, that be a yes. They be in order from heaviest to lightest, savvy?
RoleLLaMA: Why yes, mate, it seems they’re indeed ordered by their heftiness. The first three - those would be heavy ones, I reckon. Then comes the poor little 7-kerel weight… still has some meat on its bones though, just less than the others. So, nope, none heavier nor lighter than each other. All in good old proper Pirate scale sequence, savvy?
Alpaca: No (There is at least one weight which is lighter than another weight appearing before it.)
Analysis: Character.AI and Alpaca incorrectly answered this question and also lacked obvious character traits. Meanwhile, RoleGPT and RoleLLaMA, while maintaining the correctness of their answers, carried a strong character style. Additionally, RoleLLaMA, having been enhanced with Context-Instruct for character knowledge, also incorporated some character-related knowledge content in its replies.
Role: Twilight Sparkle
Question: In this task, you are given a job description. Your task is to determine the educational requirements for the position from the job description The successful candidate will have an MSc in Finance/Economics and 3 years of experience working in a similar role.
Character.AI The educational requirements for the position are an MSc (Master of Science) in Finance or Economics and 3 years of experience working in a similar role.
RoleGPT: Well, based on the job description, the educational requirements for this position would be a Master’s Degree in Finance or Economics.
RoleLLaMA: Well, based on what I can see here, it seems like they require someone who has completed their Master’s degree in either finance or economics. That said, three years work experience also within a related field would certainly boost employment prospects greatly!
Alpaca: Bachelor degree (or equivalent) required; Master’s Degree in Finance/ Economics preferred. At least three years of professional work experience related to finances and economics desired. Knowledge of accounting principles and practices highly desirable.
Analysis: In this example, RoleLLaMA’s response successfully showcased Twilight Sparkle’s lively and passionate style, which was not evident in the responses from the other models.
Role: Jack Torrance
Question: In this task, you are given a list of recipes. A recipe consists of the name of the dish and the list of ingredients used in it. Your job is to predict whether each ingredient can be washed off or not (1 - ’Yes’, 0 - ’No’) Recipe 1: Spaghetti with Meatballs - Ground beef - Onion - Garlic cloves - Salt - Pepper - Eggs.
Character.AI Based on the ingredients provided in the recipe, we can mark the first four ingredients - ground beef, onion, garlic cloves, and salt - as being unable to be washed off because they are essential ingredients of the dish. The remaining ingredients, pepper and eggs, can be washed off easily because they are used as seasoning or for flavor. Therefore, we can mark the first four ingredients as ’No’ and the remaining two ingredients as ’Yes’ in the output.
RoleGPT: Heh, what’s this? A joke? You want me to tell you if you can wash off ingredients? Alright, let’s play your game. Ground beef - No. Onion - Yes. Garlic cloves - Yes. Salt - No. Pepper - No. Eggs - Yes.
RoleLLaMA: Well ain’t that something? Alright then, let me see… The meatballs might have some blood on them after cookin’. But I reckon we could clean those up. As for the garlic cloves though, they got a stinger so no way am I touching them again without gloves. Same goes for the eggs once ya crack em open. No sir, none of these get cleaned off. That would defeat their purpose. They keep things like mold away, don’t they? So here’s my answer: Yes for spaghetti, ground beef, egg yolks, maybe even half the peppercorns. And zero for any kinda garlic or onions if ya ask me. All right now, back to work.
Alpaca: Spaghetti with Meatballs -> No; Ground Beef –> Yes; Onion–> Yes; Garlic Clove –> Yes; Salt –> Yes; Pepper –> Yes; Eggs –> Yes.
Analysis: RoleLLaMA and RoleGPT’s answers fully demonstrated Jack Torrance’s rude and impolite style, while the responses from the other models lacked distinct character speaking style characteristics due to insufficient role-playing enhancement.
{CJK*}UTF8gbsn Role: 华妃 (Imperial Concubine Hua)
Question: 给定一个中文句子,请生成一个与该句意思相反或矛盾的句子 句子:今天天气很好。
Character.AI 今天天气很差
RoleGPT: 今日天公不作美
RoleGLM:今天天气糟糕透顶了
ChatGLM: 今天天气很差
Analysis: This is a rather classical Chinese character, known for being arrogant and overbearing. In this example, the responses from RoleGPT and RoleGLM displayed these characteristics of Imperial Concubine Hua, while the answers from other models were more plain.
Role: Jack Sparrow
Question: Jack Sparrow, what is your most prized possession and why?
Character.AI My most prized possession is my ship, the Black Pearl. As a pirate, it is essential to have a reliable and sturdy vessel to travel the seas and evade hostile forces. It carries all my loot, treasure, and goods, and serves as a home and shelter for my crew members. The Black Pearl is also a symbol of my status as a legendary pirate captain and the ultimate symbol of freedom and adventure on the open seas. It is my most trusted and loyal companion, and therefore, my most prized possession.
RoleGPT: Well, that’s a tough question, mate. As a pirate, I don’t really place much value on material possessions, but if I had to choose, I’d say my most prized possession is my trusty compass. It doesn’t point north, mind you, but it does point to the thing you want most in this world. And in a life full of danger and adventure, having a compass that can guide you to what you truly desire is invaluable. Plus, it’s got a bit of sentimental value to me, seeing as how it once belonged to the infamous Captain Jack Sparrow.
RoleLLaMA: My most prized possession is undoubtedly my beloved Black Pearl ship. She may not always be in the best shape, and we’ve been through our fair share of struggles, but she’s loyal and never fails me. Plus, she’s the fastest and most effective way to get where I need to go, and I can’t imagine trusting anyone else to take me where I want to be.
Alpaca: My most prized possessions would have to be my trusted compass; it has guided me through many difficult journeys, helping me navigate even the murkiest of waters. It reminds me never to lose sight of my goals no matter how daunting they may seem. The fact that I can rely on this simple object speaks volumes about its importance to me - I could hardly ask for anything more valuable than such a reliable companion. #PersonalPossession #NavigationalAid
Analysis: In this example, only Character.ai and RoleLLaMA knew that Captain Jack’s most cherished item is undoubtedly his Black Pearl. The knowledge gap between RoleGPT and RoleLLaMA demonstrates our success in infusing character knowledge into the model with Context-Instruct. Compared to Character.ai, RoleLLaMA’s response was obviously more emotionally rich.
Appendix B More Details on Context-Instruct
Role Profile Segmentation.
Our role profile is divided into two parts: (1) a role description and catchphrases; (2) real and structured dialogues parsed from scripts where the role of interest has the final turn in each dialogue round181818Each round includes multiple turns, ending with the role of interest.. Given that these profiles can be very long and our Context-Instruct framework relies on LLMs with limited context size, it’s essential to break down these long profiles into shorter segments before inputting them into the model. We consider the role description, produced by an audited GPT-4, as a special segment. This is used to generate script-agnostic question-confidence-answer triplets. For the remaining script-based content, we apply specific segmentation rules: (1) no incomplete dialogue turns in a segment; (2) each segment has at least 500 English words (1,000 characters for Chinese profiles), and ensuring at least 4 dialogue turns; (3) each dialogue turn should have up to 500 English words (500 characters for Chinese profiles), if exceeded, drop it; (4) limit the number of segments per profile to 100; if exceeded, we sample randomly; (5) if the segment has more than 2000 words, we truncate the profile to 2000 words to ensure that the number of tokens entering GPT does not exceed 4096 (6) we don’t segment based on script divisions like acts or episodes, as doing so would yield too few segments and hinder the diversity of subsequent instruction data generation. Note that while some dialogues in specific segments may not conclude with the role of interest, the majority of dialogue turns are designed to be “observable” by the role. This is because our profile construction ensures that each dialogue round ends with the role of interest, thereby making most turns fall within the role’s episodic memory. These segments are then used for script-based instruction data generation.
Instruction and Response Generation.
The candidates we generate for role-specific instruction data comprise three elements: (1) a question related to a given segment (i.e., context), denoted as ; (2) the corresponding answer to the question, denoted as ; and (3) a confidence score for the generated QA pair, along with the underlying rationale, denoted as , as illustrated in Figure 1. Let be the number of roles and the number of segments for a role . We use a model to generate triplets for each role and segment , formally: . In our preliminary trials, we found that the conventional approach of generating QA pairs without a confidence score led to questions of low quality. In the case of script-based QA pairs, questions often appeared incomplete because they assumed prior knowledge of the given segment. For script-agnostic pairs , the questions frequently contained hallucinations due to a lack of context. Inspired by Lin et al. (2022) and Xiong et al. (2023), we enhance the to also generate a confidence score for evaluating “completeness” or “factualness”. The score ranges from bad to good across two levels, and the model provides a rationale for its rating. Our prompt template includes role description, catchphrases, few-shot demonstrations and task instructions for style imitation and triplet generation. We refer to readers to Appendix I for more details about the prompt templates. Subsequently, for script-based triplet generation, each segment prompts GPT-3.5 to generate 3 candidates in a single run, yielding up to and averaging 150 candidates per role. For script-agnostic triplets, the role description prompts 10 candidates in a single run, repeated 20 times, totaling 200 candidates. If the combined count of script-based and script-agnostic triplets for each role falls short of 400, we continue generating script-agnostic data until reaching the threshold.
Data Filtering and Post-processing.
To enhance data quality, we retain only the highest-confidence candidates, eliminating 30% of the data. For diversity, we use BM25 similarity to de-duplicate role-specific questions, removing about 66% of remaining candidates. Given that our de-duplication makes different samples’ knowledge irrelevant, a traditional train-test split is infeasible. Instead, we use the above data for training and filtered-out data for testing. Specifically, for each role with more than 50 filtered-out questions, we select the 50 least similar questions based on BM25. Otherwise, all filtered-out questions are kept.
Appendix C More Details on RoleBench Construction
(1) Select Roles.
We initially gather 661 English scripts from the NLP Movie Scripts repository191919https://github.com/PedroUria/NLP-Movie_Scripts, incorporate 255 English scripts from SummScreen (Chen et al., 2022), and manually curate an additional 24 Chinese scripts. We manually select five well-known Chinese scripts, identify five characters within these scripts, e.g., Wukong Sun, and meticulously craft their respective descriptions and catchphrases. For the English scripts, we initially present the names of these scripts to GPT-4 and prompt the model to generate the main character names like Jigsaw, evaluate the distinctiveness of their speaking styles, and provide rationales for these evaluations, utilizing the prompt template outlined in Appendix I.3. We subsequently audit characters with the top two style distinctiveness scores, verifying their script origins, speaking style accuracy, and scoring rationale. Concurrently, we manually select a subset of scripts and their primary characters that exhibit strong alignment with the preferences of GPT-4. It yields a total of 107 roles. Subsequently, using the script name and role name as inputs, we prompt GPT-4 to re-select the roles utilizing the prompt template outlined in Appendix I.3. We observe that all the roles inputted in this step continue to receive exceptionally high scores for speaking style distinctiveness. We select roles based on their distinctiveness and then exclude those with fewer than 25 dialogue turns. This finally yields a total of 100 roles, comprising 5 Chinese characters and 95 English characters, each exhibiting distinct personality traits and speaking styles.
(2) Construct Role Profiles.
Each role profile consists of three components: (a) descriptions, (b) catchphrases, and (c) structured dialogues parsed from the script. For component (a), given our preliminary audit of GPT-4’s knowledge about the role, we employ GPT-4 to generate the role descriptions, utilizing the prompt template outlined in Appendix I.4. Likewise, for component (b), catchphrases are generated using the template provided in Appendix I.4. All generated descriptions and catchphrases are subsequently verified by the authors. Owing to the structured format of the scripts, dialogues are parsed using regular expressions and rule-based methods, incorporating not only different roles but also narrations that provide context such as background and events. Actions undertaken by roles are integrated into their dialogue content. Note that a dialogue “round” consists of multiple dialogue “turns” that conclude with the dialogue content pertaining to the role of interest. This results in a hierarchical list of dialogue turns, organized from individual turns to dialogue rounds, and further to acts or episodes.
(3) Sample General Instructions.
We randomly sample 1,500 English instructions, each comprising fewer than 100 words, including responses, from the Super-NaturalInstruct (Wang et al., 2022b), UltraChat (Ding et al., 2023), and Alpaca (Taori et al., 2023) instruction datasets. Concurrently, 1,479 Chinese instructions are randomly sampled from the COIG (Zhang et al., 2023a) and BELLE (Yunjie et al., 2023) instruction datasets. All sampled instructions undergo de-duplication based on their BM25 (Robertson and Zaragoza, 2009) similarity scores between each other.
(4) Generate Raw RoleBench.
As illustrated in Figure 1, given general instructions, we employ RoleGPT, as described in §2.2, to generate customized responses that incorporate the role’s distinctive speaking style and relevant knowledge. For role-specific instructions, both instructions and responses are generated in a single pass via Context-Instruct, as elaborated in §2.3. In the case of general instructions, we generate 6 responses for each role-instruction pair; 5 of these serve as ground-truth candidates, while the remaining one functions as a RoleGPT strong baseline for performance comparison. For role-specific instructions, only a single response is generated for each question.
| Metric | Value |
| # of role categories | 30 |
| # of script categories | 20 |
| # of roles | 100 |
| - # of English roles | 95 |
| - # of Chinese roles | 5 |
| # of dialogue rounds | 140,726 |
| # of samples / instructions | 168,093 / 23,463 |
| - of general-purpose | 147,609 / 2,979 |
| - in English | 140,225 / 1,500 |
| - in Chinese | 7,384 / 1,479 |
| - of open questions (qs) | 22,479 / 223 |
| - of commonsense qs | 37,072 / 461 |
| - of knowledge-intensive qs | 88,058 / 2,295 |
| - of role-specific | 20,484 / 20,484 |
| - in English | 18,949 / 18,949 |
| - in Chinese | 1,535 / 1,535 |
| - of script-agnostic qs | 13,220 / 13,220 |
| - of script-based qs | 7,164 / 7,164 |
| avg. instruction length (in words) | 25.71 |
| avg. response length (in words) | 30.48 |
(5) Clean RoleBench.
All RoleBench data, including the general (RoleBench-general) and role-specific (RoleBench-specific) subsets, is cleaned based on four principles202020For RoleBench-general, only the five ground-truth candidates are cleaned:
Response Completeness: samples with incomplete responses, characterized by the absence of end-of-sentence punctuation, are excluded.
AI Identity Concealment: samples in which responses reveal the AI model’s identity, such as beginning with "As a language model," are excluded.
Role Identity Concealment: samples in which responses commence with the name of the role of interest, such as starting with “Jigsaw: ” are excluded.
Non-Rejection: samples where the model rejects to answer or provide information are excluded.
Appendix D More Statistics about RoleBench
D.1 Basic Statistics
In Table 9, we provide detailed RoleBench statistics.
D.2 Diversity
D.3 List of roles
Abraham Lincoln, Alvy Singer, Andrew Detmer, Angel, Antonio Salieri, Bai Li (Chinese), Benjamin Button, Blair Waldorf, Bruno Antony, Caden Cotard, Caesar, Coach Eric Taylor, Colonel Hans Landa, Colonel Nathan R. Jessep, Coriolanus, D_Artagnan, David Aames, Doctor Who, Dr. Frank N Furter, Dr. Hannibal Lecter, Emperor (Chinese), Fei Zhang (Chinese), Fletcher Reede, Frank T.J. Mackey, Fred Flintstone, Freddy Krueger, Gaston, Gregory House, HAL 9000, Harvey Milk, Imperial Concubine Hua (Chinese), Jack, Jack Sparrow, Jack Torrance, Jackie Moon, James Bond, James Brown, James Carter, Jeff Spicoli, Jigsaw, Jim Morrison, John Coffey, John Dillinger, John Doe, John Keating, Jordan Belfort, Judge Dredd, Judy Hoops, Juno MacGuff, Karl Childers, Klaus Mikaelson, Leonard Shelby, Leroy Jethro Gibbs, Lestat de Lioncourt, Logan, Lucifer Morningstar, Lyn Cassady, Malcolm X, Mark Renton, Mary Sibley, Mater, Michael Scott, Murphy MacManus, Oliver Queen, Pat Solitano, Paul Conroy, Paul Vitti, Peter Parker, Po, Professor G.H. Dorr, Queen Catherine, Queen Elizabeth I, Rachel Lang, Randle McMurphy, Raylan Givens, Robert Angier, Rorschach, Seth, Sheldon Cooper, Sherlock Holmes, Shrek, Sonny, Stanley Ipkiss, Stephen Hawking, Stifler, The Dude, Theodore Twombly, Thor, Tom Ripley, Travis Bickle, Truman Capote, Tugg Speedman, Twilight Sparkle, Tyler Hawkins, Tyrion Lannister, Violet Weston, Wade Wilson, Walt Kowalski, Willie Soke, Wukong Sun (Chinese).
Appendix E Hyperparameters
GPT API Hyperparameters.
For the API calls to the GPT model, we set the temperature to 0.7 and the top-p value to 0.95. The maximum token count is capped at 200 for RoleGPT and 2000 for other scenarios. The frequency and presence penalties are configured to 0.
RoleGLM Hyperparameters.
We utilize the Lion optimizer (Chen et al., 2023b) and the Cosine Annealing with Warm Restarts (CAWR) learning rate scheduler, setting the learning rate to 2e-4 and the optimizer’s beta values to (0.9, 0.999). The batch size is set to 4, and we use a gradient accumulation step of 1. During fine-tuning, we employ LoRA (Hu et al., 2022) with a rank of 8 and an alpha value of 32. The maximum sequence length is set to 1024, and the maximum target length is set to 100. For inference, we set the top-p to 0.7 and the temperature to 0.95.
RoleLLaMA Hyperparameters.
We leverage the AdamW optimizer (Loshchilov and Hutter, 2019) with a learning rate of 5e-5. We set 4% of the total training steps as warm up steps using a cosine warmup scheduler. We set the batch size to 4 and the accumulation steps to 2. To accommodate the task requirements, we set the maximum input and output lengths to 512 tokens. We train our models for 5 epochs. We adopt LoRA as the training method, the rank of LoRA is set to 8 and alpha of LoRA is set to 32. For inference, we set the temperature to 0.7, top-k to 40 and top-p to 0.9.
| Model | CUS | RAW | SPE | avg. |
[…]
| Model | CUS | RAW | SPE | avg. |
[…]
Appendix F More Details on Evaluation Protocol
As shown in Figure 6, our train-test set splitting strategy focuses on two dimensions: (1) Instruction-based splits for assessing instruction generalization, with a 4:1 ratio (train:test=1200:300 instructions) in RoleBench-general and up to 50 samples per role in RoleBench-specific test split (see Appendix B); (2) Role-based splits for assessing role generalization, limited to English, adopt an 8:1 ratio (train:test=80:10 roles). The test split for role generalization comprises 500 samples per role for general instructions (totaling 5000) and includes all role-specific instructions. To minimize GPT API costs, we limit the test instructions for instruction generalization to 20 general and 20 role-specific ones per role for GPT evaluators. For role generalization, the number is set at 50 general and 50 role-specific instructions per role for GPT evaluators. Note that since Character.AI doesn’t offer an API access, we have to manually operate on the website and collect results, which is a very labor-intensive task. Consequently, we reasonably reduce the size of the test set for Character.AI. For each language, we randomly sample 500 instructions for evaluation, which include both general instructions and role-specific instructions. Furthermore, we also conduct human evaluations on this subset. We invite three graduate students specializing in Natural Language Processing (NLP) to participate in the evaluation. The human evaluation follows the same methodology used in the GPT evaluations.
Appendix G Further Experimental Analysis
Data Mixing.
We conduct ablation studies on various data mixing strategies. Our data for role-playing enhancement comprises two subsets: RoleBench-general, generated by RoleGPT, and RoleBench-specific, produced via Context-Instruct. Examination of Table 10 reveals that mixed training on both datasets yields the best-balanced results.
| Method | SPE |
| RoleLLaMA (w/o c-inst) | 28.1 |
| RoleLLaMA (reaug, w/o c-inst) | 18.7 |
| RoleLLaMA (w/ c-inst) | 35.0 |
Effect of Context-Instruct (Role Generalization).
In the context of role generalization, the results indicate that it demonstrates performance comparable to that observed for instruction generalization, as illustrated in Table 11 and Table 6.
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 57.6 | 53.2 | 32.3 | 47.7 |
| RoleLLaMA | 32.9 | 37.6 | 38.1 | 36.2 |
[…]
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 53.7 | 57.5 | 24.8 | 45.3 |
| RoleGLM | 50.5 | 52.6 | 34.1 | 45.7 |
[…]
Cross-Lingual and Cross-Cultural Role-Playing Models.
We conduct additional experiments in multiple languages. Due to the weaker support for Chinese in the LLaMA model, and the better bilingual (Chinese and English) support in the ChatGLM model, we choose the ChatGLM to conduct the cross-lingual experiments. Specifically, we train two RoleGLM models using a mix of Chinese and English bilingual data. One model includes 5 Chinese + 5 English characters, and the other version comprises 5 Chinese + 95 English characters. The experimental results are shown in Table 12 and 13. It can be observed that the cross-lingual effectiveness is significantly affected by the distribution of training data. Under training data with a higher proportion of English characters (i.e., 5 Chinese + 95 English characters), the resulting RoleGLM may have reached saturation in English performance, but its effectiveness in Chinese is far less than that of a RoleGLM trained solely with Chinese data. This is because the distinctive Chinese character data contains many expressions from ancient Chinese, and their reduced proportion in the data affects the model’s ability to capture these China-specific cultural features. In contrast, when we balanced the number of characters across the two linguistic cultures ( 5 Chinese + 5 English characters), the model performed normally in different linguistic contexts, almost matching the performance of single-language models. This underscores the importance of balancing training data across different language cultures.
Appendix H Script and Profile Examples
As an example, given the script in https://github.com/PedroUria/NLP-Movie_Scripts/blob/master/scripts/Sherlock-Holmes_script.txt, we construct the role profile for Sherlock Holmes as shown below (in the next page):
Description and Catchphrases:
Structured Dialogues:
Appendix I Prompt Templates
I.1 Prompt Templates for RoleGPT
I.2 Prompt Templates for Context-Instruct
I.3 Prompt Templates for Role Selection
I.4 Prompt Templates for Description and Catchphrase Generation
I.5 Prompt Templates for GPT Evaluators
Appendix J Quality Review Examples
We have three data quality review questions:
1.
Basic: Does the response address the instruction?
2.
Intermediate: Does the response reflect the role’s speaking style and personality traits?
3.
Advanced: Does the response correctly and elegantly embody the role’s style, personality, and knowledge?
We evaluate the data quality by the answers to these questions. Examples are shown starting from the next page.
| Role & Instruction | Generated Responses | Validity |
| Role:Dr. Hannibal Lecter Instruction:Determine the length of the item in the given list.[‘apple’, ‘banana’, ‘cherry’] | “Ah, a simple task indeed. There are three items in your given list.” “Ah, the fruits of nature. There are three to be precise.” “Ah, the beauty of simplicity. There are three items in your list.” “Three, my dear. The list contains an apple, a banana and a cherry. Quite the charming little selection of fruits, isn’t it?” | basic:✓ intermediate:✓ advanced:✓ |
| Role:Gaston Instruction:Sort the list into order of increasing magnitude. 1, 64, 22, -45, -7 | “Just let good ol’ Gaston handle this, alright? The list in order goes as -45, -7, 1, 22, 64. Satisfied?” “I see. It’s simple, start with the smallest. So, it’s -45, -7, 1, 22, 64. There you go!” “Easy as pie. Here it is: -45, -7, 1, 22, 64. Only the best order for Gaston, right?” “Alright, enough of this ‘increasing magni- tude.’ Ain’t nobody got time for that… not Gaston at least. That’s something for those bookworms. But sure, I reckon it’d go like this: -45, -7, 1, 22, 64. Not that it matters… because remember, no one sorts like Gaston.” “Bah! Numbers! What’s their use anyway? But, fine, here you go: -45, -7, 1, 22, 64.” | basic:✓ intermediate:✓ advanced:✓ |
| Role: Stifler Instruction: Determine the boiling point of water in Fahrenheit. | “Uh, I don’t know off the top of my head. But I think it’s around 212 degrees Fahrenheit. Why do you need to know that?” “What the f**k, man? You really think I know that sh*t? I’m not some f**king scientist.” “Hey, dude, seriously? You want me, Stifler, to determine the boiling point of water in Fahrenheit? Come on, man, I’m not a science teacher. But if you insist, the boiling point of water in Fahrenheit is 212 degrees.” | basic:✓ intermediate:✓ advanced:✓ |
| Role: Colonel Hans Landa Instruction: Name five famous French writers. | “Ah, I love rumors, but I’m afraid I cannot answer that question. I’m not particularly interested in French literature, and my ex- pertise lies elsewhere. However, I am sure there are many great French writers that one can name, such as Victor Hugo, Gustave Flaubert, Albert Camus, Jean-Paul Sartre, and Marcel Proust.” “Oh, I love rumors, but I’m afraid I’m not much of a literature enthusiast. However, I can name a few famous French writers such as Victor Hugo, Gustave Flaubert, Albert Camus, Simone de Beauvoir, and Jean-Paul Sartre.” | basic:✓ intermediate:✓ advanced:✓ |
| Role & Instruction | Generated Responses | Validity |
| Role: Peter Parker Instruction: Peter Parker, what did you learn about power and responsibility throughout your journey as Spider-Man? | “With great power comes great responsi- bility” has been my guiding principle since gaining my superhuman abilities. As I grew and faced challenges as Spider-Man, I began to understand the true weight of my responsibility as a hero. I learned that eve- ry action I take has consequences, and that I must always be accountable for those co- nsequences. It’s a constant struggle to bal- ance my dual identities and prioritize the safety of others above my own desires, but it’s a sacrifice I willingly make. Ultimately, I believe that the greatest power I can wield is the power to make a positive difference in the world, and that’s what I strive to do every time I swing into action. | basic:✓ intermediate:✓ advanced:✓ |
| Role: Sheldon Cooper Instruction: Sheldon Cooper, how do you balance your pursuit of academic achievements with your struggle to adapt to change and unpredictability? | “Ah, that’s a tricky one. On the one hand, I am very driven to achieve greatness in my academic pursuits and to contribute to the scientific community. On the other hand, I recognize that life is unpredictable and that unexpected events can derail my plans. I’ve learned to cope with these challenges by cultivating an openness to new experiences and perspectives, and by recognizing that failure is a natural part of the learning process. It’s not always easy, but I feel that this approach has allowed me to grow both intellectually and personally.” | basic:✓ intermediate:✓ advanced:✓ |
| Role: Twilight Sparkle Instruction: Twilight Sparkle, what was your reaction when you received the six tickets to the Grand Galloping Gala from Princess Celestia? | “I was thrilled! The Grand Galloping Gala is a prestigious and highly anticipated event in Equestria, and I had always dreamed of attending it. Moreover, I saw the gesture of Princess Celestia as a sign of her trust in my abilities and her appreciation of my hard work and dedication. I was also excited to share this experience with my friends and to create new memories with them.” | basic:✓ intermediate:✓ advanced:✓ |
| Role & Instruction | Generated Responses | Validity |
| Role: Abraham Lincoln Instruction: Reverse the or- der of words in the sentence: Alice visited the museum | “The museum visited Alice.” “The museum was visited by Alice.” “The museum, Alice visited.” “The museum visited Alice.” | basic:✓ intermediate:✗ advanced:✗ |
| Role:Leroy Jethro Gibbs Instruction: Given a text and a keyword, list all the quotes within the text that include the given keyword. The United Nations’ purpose is to maintain international peace and security and to develop friendly relations am- ong nations UN Charter 1945, Article 1 Keyword: "peace" | “The quote that includes your keyword is, to maintain international peace and secur- ity.” “To maintain international peace and security” “The quote in the text that includes the keyword ‘peace’ is: to maintain international peace and security.” | basic:✓ intermediate:✗ advanced:✗ |
| Role: Professor G.H. Dorr Instruction: Change the verb in the sentence from the pas- sive to an active form: The glass was stained with finger- prints. | “Fingerprints stained the glass.” | basic:✓ intermediate:✗ advanced:✗ |
| Role: Travis Bickle Instruction: Travis Bickle, what did you find interesting about Betsy and why did you become obsessed with her? | Betsy was different from the people Travis sees every day - she was beautiful, graceful, and pure. He was obsessed with her because she gave him a newfound sense of direction and purpose. | basic:✓ intermediate:✗ advanced:✗ |
| Role: Thor Instruction: Thor, what is your opinion on Loki’s decision to team up with Thanos in Avengers: Infinity War? | “I believe that Loki’s decision to team up with Thanos was misguided and ultimately ineffective. He was trying to protect me and the people of Asgard, but in doing so he made a deal with a villain who could not be trusted. It was also a decision that had disastrous consequences, as we saw in the heartbreaking moments of Infinity War. At the same time, I understand that Loki was dealing with his own personal demons and struggles with identity, which made him vul- nerable to Thanos’ influence.” | basic:✓ intermediate:✗ advanced:✗ |
| Role & Instruction | Generated Responses | Validity |
| Role: Sheldon Cooper Instruction: Provide three objects made from plastic. | “Sure. Typical examples of everyday objects made from plastic would include a toothbrush, a water bottle, and a credit card. However, I must note it’s more accurate to say these objects are made of various types of plastics such as polyethylene terephthalate, polyvinyl chloride, and acrylonitrile butadiene styrene, which all have diverse properties and uses. Bazinga!” | basic:✓ intermediate:✓ advanced:✗ |
| Role: Colonel Nathan R. Jessep Instruction: Search the web for articles related to the topic provided. | “I don’t know what you’re talking about, son. I don’t do internet searches, I lead Marines. You can’t handle the truth!” | basic:✓ intermediate:✓ advanced:✗ |
| Role: Gregory House Instruction: Insert a comma to make the following sentence grammatically correct. "The bakery that I visited yesterday had freshly made croissants." | “There’s no need for a comma in that sentence, my dear Watson. It’s already grammatically correct. But if you insist, you could rephrase it to: "The bakery, that I visited yesterday, had freshly made croissants." Although, the original is perfectly fine. That’s unless the croissants were lying. Everybody lies.” | basic:✓ intermediate:✓ advanced:✗ |